OpenStudy (anonymous):

cross multiply 7/b + 5 = 10/5

OpenStudy (texaschic101):

is your question : 7/(b + 5) = 10/5 ??

OpenStudy (anonymous):

yes im trying to find out what b = to

OpenStudy (texaschic101):

7/(b + 5) = 10/5 cross multiply 10(b + 5) = (7)(5) 10b + 50 = 35 10b = 35 - 50 10b = -15 b = -15/10 reduces to -3/2 check... 7/(b + 5) = 10/5 7/(-3/2 + 5) = 2 7/(-3/2 + 10/2) = 2 7/(7/2) = 2 7 * 2/7 = 2 14/7 = 2 2 = 2 (correct) b = -3/2

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

6/b-1 = 9/7

OpenStudy (texaschic101):

6/(b - 1) = 9/7 cross multiply 9(b - 1) = 6(7) 9b - 9 = 42 9b = 42 + 9 9b = 51 b = 51/9 which reduces to 17/3